OpenStudy (anonymous):

Simplify the expression: (a^2b^5c^3)^4(a^3b^3)^3(a^2c)^2 a. a^13b^19c^13 b. a^21b^24c^10 c. 6(a^8b^12c^4) d. a^21b^29c^14

OpenStudy (anonymous):

its e

OpenStudy (ybarrap):

Simplify the following: (a^2 b^5 c^3)^4 (a^3 b^3)^3 (a^2 c)^2 Multiply each exponent in a^2 c by 2: (a^2 b^5 c^3)^4 (a^3 b^3)^3 a^(2×2) c^2 2×2 = 4: (a^2 b^5 c^3)^4 (a^3 b^3)^3 a^4 c^2 Multiply each exponent in a^3 b^3 by 3: (a^2 b^5 c^3)^4 a^(3×3) b^(3×3) a^4 c^2 3×3 = 9: (a^2 b^5 c^3)^4 a^(3×3) b^9 a^4 c^2 3×3 = 9: (a^2 b^5 c^3)^4 a^9 b^9 a^4 c^2 Multiply each exponent in a^2 b^5 c^3 by 4: a^(4×2) b^(4×5) c^(4×3) a^9 b^9 a^4 c^2 4×3 = 12: a^(4×2) b^(4×5) c^12 a^9 b^9 a^4 c^2 4×5 = 20: a^(4×2) b^20 c^12 a^9 b^9 a^4 c^2 4×2 = 8: a^8 b^20 c^12 a^9 b^9 a^4 c^2 a^8 b^20 c^12 a^9 b^9 a^4 c^2 = a^(8+9+4) b^(20+9) c^(12+2): a^(8+9+4) b^(20+9) c^(12+2) 12+2 = 14: a^(8+9+4) b^(20+9) c^14 20+9 = 29: a^(8+9+4) b^29 c^14 8+9+4 = 21 Answer: a^21 b^29 c^14

OpenStudy (anonymous):

wow, thanks :) That was long